# Covariant Fractons and Weitzenböck Torsion

### v1

@@ Abstract
The relation between covariant fracton gauge theory and Møller-Hayashi-Shirafuji extension of teleparallel gravity is investigated. The former is the gauge theory of a rank two symmetric tensor with gauge symmetry given by the double derivative of a scalar parameter; the latter is a class of theories which generalises Einstein gravity, whose action is quadratic in the Weitzenböck torsion. We show that the solutions of covariant fracton gauge theory can be viewed as solutions of Møller-Hayashi-Shirafuji theory, providing a new insight in the relation between covariant fractons and gravity, and elucidating the meaning of covariant fracton theories as a new type of gauge theory.
@@ 1 Introduction and outlook
In this note we pursue the investigation on covariant fracton gauge theory and its properties, studying the relation between covariant fractons and a generalisation of teleparallel gravity.
Fracton phases, which are excitations characterised by restricted mobility due to the conservation of the dipole momentum Pretko:2017xar ; Seiberg:2020bhn ; Burnell:2021reh , were firstly introduced in condensed matter physics Vijay:2015mka , but attracted interest in a broader context. In particular, covariant fracton gauge theory, introduced for the first time in Blasi:2022mbl , is the gauge theory of a rank two symmetric tensor with gauge symmetry given by the double derivative of a scalar parameter. Covariant fracton gauge theory elucidates the main properties of ordinary fracton excitations, in the context of the higher rank extensions of Maxwell theory Bertolini:2022ijb , and it can be thought as an extension of linearised Einstein gravity, invariant with respect to longitudinal diffeomorphisms Bertolini:2023juh ; Afxonidis:2023pdq . In particular, the stability of fracton gravity Afxonidis:2024tph , three-dimensional Chern-Simons-like models with covariant fracton gauge symmetry Bertolini:2024yur , and the analogue of the electromagnetic duality in covariant fracton gauge theory Bertolini:2025jov have been recently investigated.
The most general quadratic action in the derivatives of the fracton gauge field $h_{\mu\nu}^{\text{(fr)}}$ can be written in terms of the two possible scalar contractions of the gauge invariant fracton field strength Rovere:2024nwc  [^1] It appears also in Bertolini:2024apg at the linear level. $h_{\mu\nu}^{\text{(fr)}}$ and $f_{\mu\nu\varrho}^{(\text{fr})}$ are respectively denoted with $h_{\mu\nu}$ and $f_{\mu\nu\varrho}$ in Rovere:2024nwc .
@@
which is analogous to the field strength in electromagnetism. So, the action of the theory reads
$S_{\text{fr}}=\int\mathrm{d}^{d}x\,(\alpha\,{f^{\text{(fr)}}}^{\mu\nu\varrho}\,f^{\text{(fr)}}_{\mu\nu\varrho}+\beta\,{f^{\text{(fr)}}}^{\mu\nu}{}_{\nu}\,{f^{\text{(fr)}}}_{\mu\varrho}{}^{\varrho}),$ (1.2)
where $\alpha$ and $\beta$ are arbitrary constants. Taking $\alpha$ and $\beta$ such that $\alpha+\frac{1}{2}\,\beta=0$, and replacing the fracton gauge field with the parturbation of the metric in linearised gravity, one gets the Fierz-Pauli action (see (2.11)) Fierz:1939ix , which is invariant under the full group of linearised diffeomorphisms.
Since the whole dynamics of covariant fracton gauge theory is encoded in the fracton field strength, one might ask which its gravitational analogue is. We observe in this note that the covariant fracton field strength is the analogue of the linearised Weitzenböck torsion in gravity, when the latter has vanishing totally antisymmetric part. The first attempts of a torsion formulation of General Relativity, usually known as teleparallel gravity, date back to Einstein Einstein1928 , who unsuccessfully tried to unificate gravitation and electromagnetism in four dimensions, using the sixteen degrees of freedom of the vierbein, according to the works of Cartan Cartan1922 ; Cartan1923 ; Cartan1924 ; Cartan1925 and Weitzenböck Weitzenbock1928 . For the history and prehistory of the teleparallel gravity and its extension see for example Shirafuji:1995xc . In teleparallel gravity the dynamical variables of General Relativity are encoded in the torsion of a flat connection, known as Weitzenböck connection:
@@
The notion of Weitzenböck torsion is very useful for its geometrical implications, since it was employed in the geometrisation of gauge symmetries in the context of Scherk-Schwarz compactifications Scherk:1979zr , and in generalised geometry Coimbra:2011ky ; Aldazabal:2013mya ; Lee:2014mla .
@@
Møller studied extensions of teleparallel gravity, by considering the wider class of theories defined by the most general action quadratic in the Weitzenböck torsion Moller1961a ; Moller1961b ; Moller1978 . At the same time, independently, the possible phenomenological implications of the model were firstly studied by Hayashi and Shirafuji Hayashi:1979qx . The model can be also viewed as the gauge theory of translations Hayashi:1967se . The action for the theory considered by Møller and by Hayashi and Shirafuji, which we will call Møller-Hayashi-Shirafuji theory (mhs), reads
$S_{\textsc{mhs}}=\frac{1}{2}\int\mathrm{d}^{d}x\,e\,(\alpha_{1}\,T^{\mu\nu\varrho}\,T_{\mu\nu\varrho}+\alpha_{2}\,T^{\mu\nu\varrho}\,T_{\mu\varrho\nu}+\alpha_{3}\,T^{\mu\nu}{}_{\nu}\,T_{\mu\varrho}{}^{\varrho}),$ (1.4)
where $T_{\mu}{}^{\varrho}{}_{\nu}=W_{[\mu}{}^{\varrho}{}_{\nu]}$ is the Weitzenböck torsion, which is a well-defined tensor, and $\alpha_{1},\alpha_{2},\alpha_{3}$ are arbitrary constants. We will show that all the solutions of the equations of motion of covariant fracton gauge theory are solutions of the equations of motion of linearised mhs theory.
@@
To make the embedding more precise, we will take into account the brst formulation Becchi:1974md ; Becchi:1974xu ; Becchi:1975nq ; Tyutin:1975qk of covariant fracton gauge theory. In this setup, the gauge parameter is replaced by an anticommuting scalar field (fracton ghost) $\lambda$, and the gauge transformation is realised by a nilpotent differential operator $s$ on the space of the fracton gauge field and of the ghost and their derivatives, evaluated in the same spacetime point:
$s\,\lambda=0,$ (1.5a)
$s\,h_{\mu\nu}^{(\text{fr})}=\partial_{\mu}\,\partial_{\nu}\,\lambda.$ (1.5b)
The local $s$-cohomology, which classifies actions and anomalies of the theory, is generated by the undifferentiated ghost and by the fracton field strength and its derivatives, as in the analogous case of electromagnetism Rovere:2024nwc . We will show that the brst transformations of covariant fracton gauge theory can be conveniently formulated in an equivalent way, by adding two trivial doublets. The resulting transformations are
$s\,\mathring{\xi}_{\mu}=0,$ (1.6a)
$s\,\lambda=0,$ (1.6b)
$s\,h_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}+\partial_{\nu}\,\mathring{\xi}_{\mu},$ (1.6c)
$s\,b_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu},$ (1.6d)
$s\,r_{\mu}=\mathring{\xi}-\tfrac{1}{2}\,\partial_{\mu}\,\lambda,$ (1.6e)
$s\,B_{\mu\nu}=b_{\mu\nu}-\partial_{\mu}\,r_{\nu}+\partial_{\nu}\,r_{\mu},$ (1.6f)
where, in particular, the symmetric rank two tensor $h_{\mu\nu}$, which replaces the fracton gauge field, shifted by the symmetrised first derivative of $r_{\mu}$,
$h_{\mu\nu}=h_{\mu\nu}^{\text{(fr)}}+\partial_{\mu}\,r_{\nu}+\partial_{\nu}\,r_{\mu},$ (1.7)
@@
enjoys the gauge transformation of linearised gravity, given by linearised diffeomorphisms and generated by the anticommuting ghost $\mathring{\xi}$. $h_{\mu\nu}$ and $b_{\mu\nu}$, the latter being an antisymmetric rank two tensor, can be respectively viewed as the symmetric and the antisymmetric part of the linearised vielbein. In Einstein gravity, in which only $h_{\mu\nu}$ is propagating, local Lorentz transformations are included in order to gauge away $b_{\mu\nu}$. In the above transformation there is no local Lorentz transformations: this is the case of Weitzenböck geometry, which is flat and it is invariant only under global Lorentz transformations Hayashi:1977jd . Nevertheless, the last of the above transformations requires $b_{\mu\nu}$ to be cohomologically equivalent to a $\mathrm{d}$-exact two form (i.e. they differ by an $s$-exact term, which vanishes in the cohomology space). As we will see, the consequence is that the linearised Weitzenböck torsion is cohomologically equivalent to (one half) the fracton field strength, and so the totally antisymmetric part of the linearised Weitzenböck torsion is cohomologically trivial. Finally, the transformation of $r_{\mu}$ serves to implement the restriction of linearised diffeomorphisms to the “longitudinal” ones, bringing to the covariant fracton gauge transformations as they were firstly introduced Blasi:2022mbl .
@@
In Bertolini:2024apg a possible non-linear extension of covariant fracton gauge theory is discussed, by means of the Frölicher-Nijenhuis bracket, which acts on vector-valued differential forms and it coincides with the Lie derivative between vector fields when it acts on vectors FroNij1 ; FroNij2 ; FroNij3 . The non-linear extension of the fracton gauge transformation is realised by using a vector parameter, which is restricted to be the derivative of a scalar parameter only at the linear level; the gauge field is replaced by a vector-valued one-form, and a constraint is added in order to make it symmetric; a possible non-linear equation of motion is postulated, without an action from which to deduce it; the linearisation of such an equation is the equation of motion of covariant fracton gauge theory (with $\beta=0$ in (1.2)).
@@
It would be interesting to exploit the relation between covariant fractons and Weitzenböck torsion in understanding deeper this non-linear extension. In particular, it could be nice to see if it is possible to relate the postulated non-linear equation of motion with the equations of mhs theory, since it seems natural to identity the vector-valued one form with the linearised vielbein.
@@
The rest of the paper is organised as follows: in Section 2 covariant fracton theories are formulated in a way which makes manifest the relation with torsion formulation of gravity; in Section 3, 4, and 5 torsion formulation of General Relativity, the mhs theory, and its linearisation are reviewed; Section 6 is devoted to the embedding of covariant fractons in torsion gravity; finally, in Section 7 the brst formulation is investigated.
@@ 2 Covariant fracton gauge theory in “torsion-like formulation”
In this section covariant fracton gauge theory Blasi:2022mbl is formulated in a way in which it will be natural the embedding of in gravitational theories in torsion formulation.
Covariant fracton gauge theory is defined by a symmetric rank two tensor $h^{(\text{fr})}_{\mu\nu}$ and by a scalar gauge parameter, promoted to an anticommuting ghost $\lambda$ in brst formalism Becchi:1974md ; Becchi:1974xu ; Becchi:1975nq ; Tyutin:1975qk . The nilpotent brst transformations are
@@
The most general invariant quadratic action in the derivative of $h^{(\text{fr})}_{\mu\nu}$ can be written in terms of the fracton field strength, since the local brst cohomology of ghost number zero is generated by the fracton field strength and its derivatives Rovere:2024nwc . There are only two independent quadratic scalar contractions of the fracton field strength, since it transforms in the mixed Young tableau $(2,1)$. [^2] We denote with $(\lambda_{1},\dots,\lambda_{n})$ the Young tableau with $n$ rows and $\lambda_{i}$ boxes in the $i$-th row, $i=1,\dots,n$. So, the most general quadratic invariant action can be written as
$S_{\text{fr}}$ $=\int\mathrm{d}^{d}x\,(\alpha\,{f^{\text{(fr)}}}^{\mu\nu\varrho}\,f^{\text{(fr)}}_{\mu\nu\varrho}+\beta\,{f^{\text{(fr)}}}^{\mu\nu}{}_{\nu}\,{f^{\text{(fr)}}}_{\mu\varrho}{}^{\varrho})=$
@@
$E^{\text{(fr)}}_{\mu\nu}:=-\frac{1}{2}\,\frac{\delta\,S_{\text{fr}}}{\delta\,{h^{(\text{fr})}}^{\mu\nu}}=\alpha\,\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho}{}_{(\mu\nu)}-\tfrac{\beta}{2}\,\partial_{(\mu}\,{f^{\text{(fr)}}}_{\nu)\varrho}{}^{\varrho}+\beta\,\eta_{\mu\nu}\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho\sigma}{}_{\sigma},$ (2.5)
@@
Indeed, if one replace $h^{\text{(fr)}}_{\mu\nu}$ in the action (2) with a rank two symmetric tensor $h_{\mu\nu}$, thought as the perturbation of a metric $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$, and $\alpha$ and $\beta$ are such that
$\alpha+\tfrac{\beta}{2}=0,$ (2.10)
choosing the normalisation $\alpha=\frac{1}{2}$, one recovers the Fierz-Pauli action Fierz:1939ix
$S_{\textsc{fp}}=\tfrac{1}{2}\int\mathrm{d}^{d}x\,(\partial_{\varrho}\,h_{\mu\nu}\,\partial^{\varrho}\,h^{\mu\nu}-2\,\partial^{\nu}\,h_{\mu\nu}\,\partial_{\varrho}\,h^{\mu\varrho}+2\,\partial_{\mu}\,h_{\nu}{}^{\nu}\,\partial_{\varrho}\,h^{\mu\varrho}-\partial_{\mu}\,h_{\nu}{}^{\nu}\,\partial^{\mu}\,h_{\varrho}{}^{\varrho}).$ (2.11)

### v3

@@ Abstract
The relation between covariant fracton gauge theory and Møller-Hayashi-Shirafuji theory of gravity is investigated. The former is the gauge theory of a rank-two symmetric tensor with gauge symmetry given by the double derivative of a scalar parameter; the latter is the most general theory, whose action is quadratic in the Weitzenböck torsion. We show that the solutions of covariant fracton gauge theory describe a subsector of the space of solutions of Møller-Hayashi-Shirafuji theory, providing a new insight in the relation between covariant fractons and gravity, and elucidating the meaning of covariant fracton theory as a new type of gauge theory.
@@ 1 Introduction and outlook
In this paper we pursue the investigation on covariant fracton gauge theory and its properties, studying the relation between covariant fractons and teleparallel gravity.
Fracton phases, which are excitations characterised by restricted mobility due to the conservation of the dipole moment Pretko:2017xar ; Seiberg:2020bhn ; Burnell:2021reh , were firstly introduced in condensed matter physics Vijay:2015mka , but attracted interest in a broader context. In particular, covariant fracton gauge theory, introduced for the first time in Blasi:2022mbl , is the gauge theory of a rank-two symmetric tensor $h_{\mu\nu}^{(\text{fr})}$ with gauge symmetry given by the double derivative of a scalar parameter
$\delta\,h^{\text{(fr)}}_{\mu\nu}=\partial_{\mu}\,\partial_{\nu}\,\lambda.$ (1.1)
Covariant fracton gauge theory elucidates the main properties of ordinary fracton excitations, in the context of the higher rank extensions of Maxwell theory Bertolini:2022ijb , and it can be thought as an extension of linearised Einstein gravity, invariant with respect to longitudinal diffeomorphisms Bertolini:2023juh ; Afxonidis:2023pdq . Recently, three-dimensional Chern-Simons-like models with covariant fracton gauge symmetry Bertolini:2024yur , and the analogue of the electromagnetic duality in covariant fracton gauge theory Bertolini:2025jov have been investigated.
[…]
@@
which is analogous to the field strength in electromagnetism. The action reads
[…]
where $\alpha$ and $\beta$ are arbitrary constants. Taking $\alpha$ and $\beta$ such that $\frac{\beta}{\alpha}=-2$, and replacing the fracton gauge field with the perturbation of the metric in linearised gravity, one gets the Fierz-Pauli action (see equation (2.12)) Fierz:1939ix , which is invariant under the full group of linearised diffeomorphisms.
Since the whole dynamics of covariant fracton gauge theory is encoded in the fracton field strength, one might ask which its gravitational analogue is. We observe in this paper that the covariant fracton field strength is the analogue of the linearised Weitzenböck torsion in gravity, when the linearised vielbein has vanishing totally antisymmetric part. The first attempts of a torsion formulation of General Relativity, usually known as teleparallel equivalent of General Relativity, date back to Einstein Einstein1928 , who unsuccessfully tried to unificate gravitation and electromagnetism in four dimensions, using the sixteen degrees of freedom of the vierbein, according to Cartan’s Cartan1922 ; Cartan1923 ; Cartan1924 ; Cartan1925 and Weitzenböck’s Weitzenbock1928 works. For the history and prehistory of the teleparallel gravity see for example Shirafuji:1995xc ; Maluf:2013gaa . In teleparallel gravity the dynamical variables are encoded in the torsion of a flat connection, known as Weitzenböck connection:
@@
where $e_{\mu}{}^{a}$ is the vielbein, and $e^{\mu}{}_{a}$ its inverse. The Weitzenböck torsion, which is the antisymmetric part of the previous connection
$T_{\mu}{}^{\varrho}{}_{\nu}=W_{[\mu}{}^{\varrho}{}_{\nu]}=e^{\varrho}{}_{a}\,\partial_{[\mu}\,e_{\nu]}{}^{a}$, (1.5)
@@
is a well-defined covariant tensor, which is very useful for its geometrical implications: namely, it was employed in the geometrisation of gauge symmetries in the context of Scherk-Schwarz compactifications Scherk:1979zr , and in generalised geometry Coimbra:2011ky ; Aldazabal:2013mya ; Lee:2014mla . Møller studied extensions of General Relativity in teleparallel formulation, by considering the wider class of theories defined by the most general action quadratic in the Weitzenböck torsion Moller1961a ; Moller1961b ; Moller1978 . At the same time, independently, the possible phenomenological implications of the model were first studied by Hayashi and Shirafuji Hayashi:1979qx . The model can be also viewed as the gauge theory of translations Hayashi:1967se . The action for the theory considered by Møller and by Hayashi and Shirafuji, which we will call Møller-Hayashi-Shirafuji theory (mhs), reads
[…]
where $\alpha_{1},\alpha_{2},\alpha_{3}$ are arbitrary constants.
We will show that, when $\frac{\beta}{\alpha}\neq 2$, the solutions of the equations of motion of covariant fracton gauge theory can be viewed as the subsector in the space of solutions of the equations of motion of linearised mhs theory, identified by the vanishing antisymmetric part of the linearised vielbein, denoted with $b_{\mu\nu}$,
$b_{\mu\nu}=0.$ (1.7)
The embedding of covariant fracton solutions in mhs theory is realised when the fracton gauge field $h_{\mu\nu}^{(\text{fr})}$ is identified with the symmetric part of the linearised vielbein, denoted with $h_{\mu\nu}$,
$h_{\mu\nu}^{\text{(fr)}}=h_{\mu\nu},$ (1.8)
and the constants $\alpha$ and $\beta$ in the covariant fracton action (1.3) are identified with combinations of the constant $\alpha_{1},\alpha_{2}$, and $\alpha_{3}$ in the mhs action (1.6)
$\alpha=\tfrac{1}{4}\,(2\,\alpha_{1}+\alpha_{2}),\quad\beta=\tfrac{1}{2}\,\alpha_{3}.$ (1.9)
To show this, focusing on four dimensions for definiteness, we will compare the particle spectrum of both theories. Except for some particular cases, tuned by special values of the constants $\alpha_{1},\alpha_{2}$, and $\alpha_{3}$, such that the gauge invariance is enlarged, the spectrum of mhs theory in four dimensions consists in six degrees of freedom, given by a particle with helicity $\pm 2$, one with helicity $\pm 1$, and two scalar particles. The antisymmetric field $b_{\mu\nu}$ describes only a scalar particle (as expected, since a two-form gauge field in four dimensions is dual to a scalar field). The sector with vanishing $b_{\mu\nu}$ coincides with the particle spectrum of covariant fracton gauge theory, which indeed describes the remaining five propagating degrees of freedom Afxonidis:2023pdq , except for the particular cases in which the gauge symmetry is enlarged.
Linearised diffeomorphisms act on $b_{\mu\nu}$ as a one-form gauge transformation for a two-form field (usually known as Ramond-Kalb field Kalb:1974yc ). The condition (1.7) is not diffeomorphism invariant, but it leaves a residual gauge symmetry when the gauge parameter is exact, that is, when it is a derivative. This is the case of longitudinal diffeomorphisms, so that the residual gauge symmetry is the fracton gauge symmetry. In a sense, covariant fracton solutions spontaneously break the diffeomorphism symmetry of mhs theory, preserving only the fracton gauge symmetry.
@@
To elucidate this setting, we will take into account the brst formulation Becchi:1974md ; Becchi:1974xu ; Becchi:1975nq ; Tyutin:1975qk of covariant fracton gauge theory. In the brst setup, the gauge parameter is replaced by an anticommuting scalar field (fracton ghost) $\lambda$, and the gauge transformation is realised by a nilpotent differential operator $s$ on the space of the fracton gauge field and of the ghost and their derivatives, evaluated in the same spacetime point. The local $s$-cohomology, which classifies the consistent anomalies, is generated by the undifferentiated ghost and by the fracton field strength and its derivatives, as in the analogous case of electromagnetism Rovere:2024nwc .
@@
We will show that the brst transformations of covariant fracton gauge theory can be conveniently formulated in an equivalent way, by including two brst trivial doublets. In this formulation a symmetric rank-two tensor $h_{\mu\nu}$ and an antisymmetric one $b_{\mu\nu}$ are included. They can be respectively viewed as the symmetric and the antisymmetric part of the linearised vielbein, but the brst transformations do not involve any local Lorentz transformation: this is the case of Weitzenböck geometry, which is flat and it is invariant only under global Lorentz transformations Hayashi:1977jd . Nevertheless, on the one side, the brst transformations require the restriction of linearised diffeomorphisms to the “longitudinal” ones, bringing us to covariant fracton gauge transformations as they were first introduced Blasi:2022mbl ; on the other side, $b_{\mu\nu}$ is equivalent to zero. This individuates the subsector of mhs theory in which covariant fracton gauge theory lives.
Covariant fractons were introduced in Blasi:2022mbl only at linear level. In Afxonidis:2024tph it is shown that the theory is afflicted by instabilities, which in principle can be cured by considering suitable interactions. Going towards this direction, the result in this note can be viewed as a possible way to embed covariant fracton theory in a larger theory of gravity, whose non-linear extension is automatically given by the extension of General Relativity in teleparallel formulation, studied by Møller, Hayashi, and Shirafuji.
@@
A possible non-linear extension of covariant fracton gauge theory is also discussed in Bertolini:2024apg , where the Frölicher-Nijenhuis bracket is employed, acting on vector-valued differential forms and coinciding with the Lie derivative between vector fields when it acts on vectors FroNij1 ; FroNij2 ; FroNij3 . The non-linear extension of the fracton gauge transformation is realised by using a vector parameter, which is restricted to be the derivative of a scalar parameter only at the linear level; the gauge field is replaced by a vector-valued one-form, and a constraint is added in order to make it symmetric; a possible non-linear equation of motion is postulated, without an action from which to deduce it; the linearisation of such an equation is the equation of motion of covariant fracton gauge theory (with $\beta=0$ in (1.3)).
@@
It would be interesting to exploit the relation between covariant fractons and Weitzenböck torsion in understanding deeper this non-linear extension. In particular, one could try to see if it is possible to relate the postulated non-linear equation of motion with the equations of mhs theory, since it seems natural to identity the vector-valued one form with the linearised vielbein.
@@
The rest of the paper is organised as follows: in Section 2 covariant fracton theory is formulated in a way which makes manifest the relation with torsion formulation of gravity; in Section 3, 4, and 5 torsion formulation of General Relativity, the mhs theory, and its linearisation are reviewed; Section 6 and 7 the particle spectrum of covariant fracton and mhs theory are computed and compared, providing the embedding of covariant fractons in mhs theory; finally, in Section 8 the equivalent brst formulation is investigated. In Appendix A there is a summary of Weitzenböck geometry; in Appendix B a computation for brst cohomology is shown; Appendix C is devoted to a comparison between different conventions; and in Appendix D there is a summary of polarisation matrices.
@@ 2 Covariant fracton gauge theory in “torsion-like formulation”
In this section covariant fracton gauge theory Blasi:2022mbl is formulated in a way in which its embedding in gravitational theories in torsion formulation will be natural.
Covariant fracton gauge theory is defined by a symmetric rank-two tensor $h^{(\text{fr})}_{\mu\nu}$ and by a scalar gauge parameter, promoted to an anticommuting ghost $\lambda$ in brst formalism Becchi:1974md ; Becchi:1974xu ; Becchi:1975nq ; Tyutin:1975qk . The nilpotent brst transformations are
@@
The most general invariant quadratic action in the derivative of $h^{(\text{fr})}_{\mu\nu}$ can be written in terms of the fracton field strength. There are only two independent quadratic scalar contractions of the fracton field strength, since it transforms in the mixed Young tableau (2,1)(2,1). [^2] We denote with $(\lambda_{1},\dots,\lambda_{n})$ the Young tableau with $n$ rows and $\lambda_{i}$ boxes in the $i$th row, $i=1,\dots,n$. $(\lambda_{1},\dots,\lambda_{n})_{t}$ is the corresponding traceless part. So, the most general quadratic invariant action can be written as
[…]
@@
$E^{\text{(fr)}}_{\mu\nu}:=-\frac{1}{2}\,\frac{\delta\,S_{\text{fr}}}{\delta\,{h^{(\text{fr})}}^{\mu\nu}}$ $=\alpha\,\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho}{}_{(\mu\nu)}-\tfrac{\beta}{2}\,\partial_{(\mu}\,{f^{\text{(fr)}}}_{\nu)\varrho}{}^{\varrho}+\beta\,\eta_{\mu\nu}\,\partial_{\varrho}\,{f^{\text{(fr)}}}^{\varrho\sigma}{}_{\sigma}$= (2.5)
$=2\,\alpha\,\partial^{2}\,h_{\mu\nu}^{\text{(fr)}}-\tfrac{2\,\alpha-\beta}{2}\,\partial_{(\mu}\,\partial^{\varrho}\,h^{\text{(fr)}}_{\nu)\varrho}-\beta\,\partial_{\mu}\,\partial_{\nu}\,h^{\text{(fr)}\,\varrho}{}_{\varrho}\,$+
$\quad+\beta\,\eta_{\mu\nu}\,(\partial^{2}\,h^{\text{(fr)}\,\varrho}{}_{\varrho}-\partial_{\varrho}\,\partial_{\sigma}\,h^{\text{(fr)}\,\varrho\sigma})$ (2.6)
@@
Indeed, if one replace $h^{\text{(fr)}}_{\mu\nu}$ in the action (2) with a rank-two symmetric tensor $h_{\mu\nu}$, thought as the perturbation of a metric $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$, and $\alpha$ and $\beta$ are such that
$\tfrac{\beta}{\alpha}=-2,$ (2.11)
choosing the normalisation $\alpha=\frac{1}{2}$, one recovers the massless Fierz-Pauli action Fierz:1939ix
[…]